OpenStudy (anonymous):

Find any critical numbers of the function. f(θ) = 10sec θ + 5tan θ, 0 < θ < 2π

OpenStudy (anonymous):

HEY

OpenStudy (anonymous):

10secxtanx+5sec^2x=0 2tanx=-secx 2sinx=-1 sinx=-1/2 sin^-1(-1/2)=x x=-pi/6 or 11pi/6

OpenStudy (anonymous):

r u sure thats right cause i put it in and it marked it wrong

OpenStudy (zarkon):

\[\frac{7\pi}{6},\frac{11\pi}{6}\]